# 2nd Workshop on Multi-modal Foundation Models and Large Language Models for Life Sciences

### v1

@@ Abstract
In-context learners like TabPFN are promising for biomolecule efficacy prediction, where established molecular feature sets and relevant experimental results can serve as powerful contextual examples. However, their performance is highly sensitive to the provided context, making strategies like post-hoc ensembling of models trained on different data subsets a viable approach. An open question is how to select the best models for the ensemble without access to ground truth labels. In this study, we investigate an uncertainty-guided strategy for model selection. We demonstrate on an siRNA knockdown efficacy task that a TabPFN model using simple sequence-based features can surpass specialized state-of-the-art predictors. We also show that the model’s predicted inter-quantile range (IQR), a measure of its uncertainty, has a negative correlation with true prediction error. By selecting and averaging an ensemble of models with the lowest mean IQR, we achieve superior performance compared to naive ensembling or using a single model trained on all available data. This finding highlights model uncertainty as a powerful, label-free heuristic for optimizing biomolecule efficacy predictions.
@@ 1 Introduction
Strategies to address this include data retrieval and fine-tuning, as exemplified by LoCalPFN [7], or post-hoc ensembling [8], which combines predictions from models trained on different data subsets. However, effective ensembling still requires a reliable method for selecting best performing models. In this paper, we explored an uncertainty based strategy for this task. We propose that a model’s predicted uncertainty can serve as a proxy for its true prediction error, allowing us to select and combine models with the lowest predicted uncertainty to form a final, more accurate ensemble. While prior work has used uncertainty of sensitive attributes to guide sample selection to improve fairness in ICLs [9], its use for improving model accuracy by model selection has not been explored to the best of our knowledge. We demonstrated on an siRNA knockdown efficacy task that TabPFN model with simple features surpasses specialized, state-of-the-art models. Furthermore, we show the model’s inter-quantile range (IQR) reflects its prediction uncertainty and negatively correlates with true prediction error, making it a competitive metric for selecting models that yield optimal performance.
@@ 3.3 Model Selection with Inter-Quantile Range
Post-hoc ensembling (PHE) is a common technique to improve model performance by averaging predictions from multiple models trained with different model hyperparameters and different datasets. Given that IQR is indicative of model performance, we investigated its use as a metric for model selection within a PHE framework. For each test set, we trained an ensemble of 400 models, where each model was trained on $k$ randomly selected training subsets ($k$ was chosen randomly between 1 and 20 for each model). The process was repeated 3 times using different random seeds. We then compared three strategies: (1) a single baseline model trained on all available training data (All data single model), (2) a naive ensemble that averages predictions from all 400 models (Full ensemble mean), and (3) our uncertainty-guided ensemble that averages predictions from only the top 10% of models with the lowest mean IQR (Top 10% ensemble mean). All results are given as means and 95% confidence intervals of 3 parallel runs.
@@
Dataset Top 10% Full ensemble All data Ensemble best
ensemble mean mean single model (oracle result)
@@
Dataset Top 10% Full ensemble All data Ensemble best
ensemble mean mean single model (oracle result)
@@
The results are summarized in Table 2 (MAE) and Table 3 (Correlation). While our selection strategy has a marginal effect on MAE, it provides a notable improvement in correlation for two of the three test datasets. For Target1 (A), the top-10% ensemble yields a correlation of 0.278 $\pm$ 0.015, a substantial improvement over the performance of a single model trained with all data (correlation of 0.051). Similarly, for Target2, the top-10% ensemble achieves the highest correlation, highlighting the effectiveness of our uncertainty-guided approach. While Target1 (B) remains a challenging case for all methods, these results demonstrate that IQR-based model selection is a promising strategy. We would like to emphasize that the extra amount of compute is acceptable, given that each model in the ensemble only operates on a limited size of data, and the inference with different models in the ensemble is embarrassingly parallelizable. Furthermore, this approach provides a direct way to handle large amount of context data that do not fit into a single forward pass of the TabPFN model. The "Ensemble Best" column shows the performance of the single best model in the ensemble from all 3 parallel runs. It is an oracle result which means there is no way to identify a single best model a priori for new test data. Comparing our approach with the "ensemble best" reveals the upper bound of performance and the remaining gap for improvement in model selection strategies.
@@ 5 Conclusion
We have shown that in-context learners like TabPFN, equipped with straightforward engineered features, could surpass specialized state-of-the-art models on siRNA knockdown efficacy prediction. Our primary contribution is the finding that the model’s self-reported uncertainty, quantified by the inter-quantile range (IQR) of its predictions, is a potent, label-free heuristic for post-hoc model selection. We showed that an ensemble composed of models selected for their low prediction uncertainty can achieve superior performance, particularly in terms of correlation, compared to a single model using all available data or a naive ensemble average. This approach offers a practical strategy to enhance the reliability of in-context learners in real-world scenarios where ground truth labels are unavailable for model tuning. Future work should aim to narrow the gap to the oracle "ensemble best" by further iterating on the algorithm, and by validating this uncertainty-guided methodology across a broader range of biomolecule prediction tasks.

### v2

@@ Abstract
In-context learners like TabPFN are promising for biomolecule efficacy prediction, where established molecular feature sets and relevant experimental results can serve as powerful contextual examples. However, their performance is highly sensitive to the provided context, making strategies like post-hoc ensembling of models trained on different data subsets a viable approach. An open question is how to select the best models for the ensemble without access to ground truth labels. In this study, we investigate an uncertainty-guided strategy for model selection. We demonstrate on an siRNA knockdown efficacy task that a TabPFN model using straightforward sequence-based features can surpass specialized state-of-the-art predictors. We also show that the model’s predicted inter-quantile range (IQR), a measure of its uncertainty, has a negative correlation with true prediction error. We developed the OligoICP method, which selects and averages an ensemble of models with the lowest mean IQR for siRNA efficacy prediction, achieving superior performance compared to naive ensembling or using a single model trained on all available data. This finding highlights model uncertainty as a powerful, label-free heuristic for optimizing biomolecule efficacy predictions.
@@ 1 Introduction
Strategies to address this include data retrieval and fine-tuning, as exemplified by LoCalPFN [7], or post-hoc ensembling (PHE)[8], which combines predictions from models trained on different data subsets. However, effective ensembling still requires a reliable method for selecting best performing models. In this paper, we introduce OligoICP (Oligo In-Context Predictor), an uncertainty based strategy for model selection in PHE for siRNA knockdown efficacy prediction. The core principle of OligoICP is that a model’s predicted uncertainty can serve as a proxy for its true prediction error, allowing us to select and combine models with the lowest predicted uncertainty to form a final, more accurate ensemble. While prior work has used uncertainty of sensitive attributes to guide sample selection to improve fairness in ICLs [9], its use for improving model accuracy by model selection has not been explored to the best of our knowledge. We demonstrated that our method, based on TabPFN model with straightforward features, surpasses specialized, state-of-the-art models. Furthermore, we show the model’s inter-quantile range (IQR) reflects its prediction uncertainty and negatively correlates with true prediction error, making it a competitive metric for selecting models that yield optimal performance.
@@ 3.3 Model Selection with Inter-Quantile Range
Post-hoc ensembling (PHE) is a common technique to improve model performance by averaging predictions from multiple models trained with different model hyperparameters and different datasets. Given that IQR is indicative of model performance, we developed the OligoICP framework, which uses IQR as a metric for model selection. For each test set, we trained an ensemble of 400 models, where each model was trained on $k$ randomly selected training subsets ($k$ was chosen randomly between 1 and 20 for each model). The process was repeated 3 times using different random seeds. We then compared three strategies: (1) a single baseline model trained on all available training data (All data single model), (2) a naive ensemble that averages predictions from all 400 models (Full ensemble mean), and (3) OligoICP, which averages predictions from only the top 10% of models with the lowest mean IQR (Top 10% ensemble mean). All results are given as means and 95% confidence intervals of 3 parallel runs.
@@
Dataset OligoICP (Top 10% Full ensemble All data Ensemble best
ensemble mean) mean single model (oracle result)
@@
Dataset OligoICP (Top 10% Full ensemble All data Ensemble best
ensemble mean) mean single model (oracle result)
@@
The results are summarized in Table 2 (MAE) and Table 3 (Correlation). While the OligoICP approach has a marginal effect on MAE, it provides a notable improvement in correlation for two of the three test datasets. For Target1 (A), OligoICP yields a correlation of 0.278 $\pm$ 0.015, a substantial improvement over the performance of a single model trained with all data (correlation of 0.051). Similarly, for Target2, OligoICP achieves the highest correlation, highlighting the effectiveness of our ensemble selection approach. While Target1 (B) remains a challenging case for all methods, these results demonstrate that IQR-based model selection is a promising strategy. We would like to emphasize that the extra amount of compute is acceptable, given that each model in the ensemble only operates on a limited size of data, and the inference with different models in the ensemble is embarrassingly parallelizable. Furthermore, OligoICP provides a direct way to handle large amount of context data that do not fit into a single forward pass of the TabPFN model. The "Ensemble Best" column shows the performance of the single best model in the ensemble from all 3 parallel runs. It is an oracle result which means there is no way to identify a single best model a priori for new test data. Comparing our approach with the "ensemble best" reveals the upper bound of performance and the remaining gap for improvement in model selection strategies.
@@ 5 Conclusion
We have shown that in-context learners like TabPFN, equipped with straightforward engineered features, could surpass specialized state-of-the-art models on siRNA knockdown efficacy prediction. Our primary contribution is the development and validation of OligoICP, a method that uses the model’s self-reported uncertainty, quantified by the inter-quantile range (IQR) of its predictions as a potent, label-free heuristic for post-hoc model selection. We showed that the OligoICP ensemble, composed of models selected for their low prediction uncertainty, can achieve superior performance, particularly in terms of correlation, compared to a single model using all available data or a naive ensemble average. This approach offers a practical strategy to enhance the reliability of in-context learners in real-world scenarios where ground truth labels are unavailable for model tuning. Future work should aim to narrow the gap to the oracle "ensemble best" by further iterating on the algorithm, and by validating this uncertainty-guided methodology across a broader range of biomolecule prediction tasks.